Cathy and Tom's Specialty Ice Cream Company operates a small production facility for the local community. The facility has the capacity to make 25,200 gallons of the single flavor, GUI Chewy, annually. The plant has only two customers, Chuck's Gas & Go and Marcee's Drive & Chew DriveThru. Annual orders for Chuck's total 12,600 gallons and annual orders for Marcee's total 6,300 gallons. Variable manufacturing costs are $1.20 per gallon, and annual fixed manufacturing costs are $31,500.

The ice cream business has two seasons, summer and winter. Each season lasts exactly six months. Chuck's orders 6,300 gallons in the summer and 6,300 gallons in the winter. Marcee's is closed in the winter and orders all 6,300 gallons in the summer.

Required:

a. Calculate the product cost for each season with excess capacity costs assigned to season in which it is incurred. (Round your intermediate calculations and final answers to 2 decimal places.)

winter per gallon

summer per gallon

b. Calculate the product cost for each season with excess capacity costs assigned to the season requiring it. (Round your intermediate calculations and final answers to 2 decimal places.)

winter per gallon

summer per gallon
Answer #1

SOLUTION

A. Capacity cost in each season = 31,500 /2 = 15,750

Winter-

Overhead rate = 15,750 / 6,300 gallons = $2.50

Production cost = ($1.20 + $2.50) = $3.70 per gallon

Summer-

Overhead rate = 15,750 /12,600 gallons = $1.25

Production cost = ($1.20 + $1.25) = $2.45 per gallon

B.

Winter-

Overhead rate = (15,750*50%) / 6,300 gallons = $1.25

Production cost = ($1.20 + $1.25) = $2.45 per gallon

Summer-

Overhead rate = (15,750 + (15,750*50%)) /12,600 gallons = $1.88

Production cost = ($1.20 + $1.88) = $3.08 per gallon